OpenStudy (breanna1999):

Given the parent functions f(x) = log10 x and g(x) = 3x − 1, what is f(x) • g(x)? f(x) • g(x) = log10 (3x − 1)x f(x) • g(x) = log10 x3x − 1 f(x) • g(x) = 3x log10 x + log10 x f(x) • g(x) = log10 x − 3x log10 x

OpenStudy (therawbugeyes):

that is C i belive

OpenStudy (breanna1999):

i thought it was A

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{f=(\log x)} \qquad\qquad\qquad \color{orangered}{g=(3x-1)}\]Then,\[\large\rm \color{royalblue}{f}\cdot\color{orangered}{g}\quad=\quad \color{royalblue}{(\log x)}\cdot\color{orangered}{(3x-1)}\]

satellite73 (satellite73):

put a times sign between them

zepdrix (zepdrix):

Distribute the log x to each term.